OpenStudy (anonymous):

I'm having issues with figuring out set-builder notation, I have a general idea of how it works, but this is giving me problems. 4y + 7 (the symbol is greater than or equal to) >_ 23

OpenStudy (freckles):

well first do you know how to solve the inequality 4y+7>=23?

OpenStudy (anonymous):

I've been watching some tutorials on it, but its kind of going over my head.... math is not my best subject.... I'm pretty sure its starting with {y|(its this bit that's giving me issues) }

OpenStudy (freckles):

well to get to that part you need to first isolate y in the inequality 4y+7>=23

OpenStudy (freckles):

I would start by undoing that addition by 7 there

OpenStudy (freckles):

to undo addition you subtract whatever you do to one side you do to the other so 4y+7>=23 so try subtracting 7 on both sides as a first step

OpenStudy (freckles):

what do you after subtracting 7 on both sides?

OpenStudy (anonymous):

So its going something like: 4y + 7 - 7 > 23 - 7?

OpenStudy (anonymous):

yes?

OpenStudy (freckles):

right

OpenStudy (freckles):

and the whole reason we did that is so we could isolate the 4y

OpenStudy (freckles):

\[4y \ge 16 \]

OpenStudy (freckles):

now to undo the multiplication by 4 you need to divide by 4

OpenStudy (freckles):

whatever you do to one side do to the other

OpenStudy (freckles):

the inequality sign only changes direction when you multiply or divide both sides by a negative number so the inequality keeps its direction is is already in

OpenStudy (freckles):

so try to divide both sides by 4 and let me know what you have as a result

OpenStudy (freckles):

\[\left\{ y: y \ge 4 \right\}\]

OpenStudy (freckles):

you were right I thought

OpenStudy (anonymous):

Ah, that make so much more sense now. Ty

OpenStudy (anonymous):

Thanks for explaining it!

OpenStudy (freckles):

np

OpenStudy (freckles):

were you trying to skip the solving part and go straight to building your set?

OpenStudy (anonymous):

No, I needed an understanding of how it works. I like to see an example of how to solve, it gives me an idea for how to do other problems.

OpenStudy (anonymous):

Now I can actually move on in progress on my assignment. That was an example problem form my book. The book explains in it very techinical terms that are harder for me to understand

OpenStudy (anonymous):

*from

OpenStudy (freckles):

here is another example and I wanted to include this one just in case you end up multiplying and dividing both sides by a negative \[-4x-1 \le 6 \\ \text{ add one \to both sides } \\ -4x \le 6+1 \\ -4x \le 7 \\ \text{ divide both sides by -4 } \\ \text{ remember when dividing both sides } \\ \text{ or multiplying both sides by a \neg flip the inequality direction } \\ x \ge \frac{7}{-4} \\ \left\{ x: x \ge \frac{-7}{4} \right\}\]

OpenStudy (anonymous):

So the < flips when the problem has division? And > flips when there is multiplication?

OpenStudy (freckles):

when you have division or multiplication of a negative number on both sides for example we know that 3>2 but if you multiply both sides by -1 then you -3<-2

OpenStudy (freckles):

notice -3>-2 isn't true at all

OpenStudy (freckles):

but -3<-2 is true

OpenStudy (freckles):

so if you have -x>1 then x<-1

OpenStudy (anonymous):

So, its changing the < or > to make the equation true

OpenStudy (freckles):

yes

OpenStudy (freckles):

only when dividing and/or multiplying both sides by a NEGATIVE number

OpenStudy (anonymous):

Ok, that helps a lot, ty

OpenStudy (freckles):

like if you have 5x>10 then x>2

OpenStudy (freckles):

but if -5x>10 then x<-2

OpenStudy (freckles):

of if -5x>-10 then x<2 (<--one more example)

OpenStudy (anonymous):

Awesome, that makes sense to me now, thank you very very much, you have been a great help! =)